# Modeling and Beamforming Optimization for Pinching-Antenna Systems

### v2

@@ IV-B Transmit Power Versus the Minimum SINR
IV-C Transmit Power Versus the Distance
@@
IV-D Transmit Power Versus the Number of Antennas
Figure 10: Transmit power versus the number of antennas.
@@ IV-D Transmit Power Versus the Number of Antennas
IV-E Transmit Power Versus the Number of Discrete Positions
@@ V Conclusions
However, several key challenges remain unresolved. For instance, the proposed physics model has been applied only to the downlink scenario, whereas modeling for uplink scenarios may be more complex and requires further investigation. Additionally, this study considered the simplest transmit beamforming architecture, where each waveguide is connected to a single RF chain. Exploring hybrid analog-digital beamforming architectures with pinching beamforming, which support fully connected configurations and a flexible number of RF chains, is another promising direction for future research.

### v3

@@ IV-B Impact of the Minimum SINR
IV-C Impact of the Distance
Figure 8: Transmit power versus the minimum SINR.
@@ IV-C Impact of the Distance
Figure 10: Transmit power versus the number of antennas.
@@
IV-D Impact of the Number of Antennas
@@
Figure 12: Average SINR versus the channel error bound $\varepsilon_{\mathrm{est}}$.
IV-E Impact of the Number of Discrete Positions
@@
IV-F Impact of the Channel Estimation Error
Fig. 12 explores the impact of the channel estimation error on the average SINR among users. We adopt a deterministic model to characterize the channel estimation error, which models the free-space channel as [35]
$\mathbf{h}_{k}=\hat{\mathbf{h}}_{k}+\Delta\mathbf{h}_{k},\quad\|\Delta\mathbf{h}_{k}\|\leq\varepsilon_{\mathrm{est}},\forall k,$ (71)
where $\hat{\mathbf{h}}_{k}$ is the estimated channel for user $k$ and $\Delta\mathbf{h}_{k}$ denotes the channel estimation error with its norm bounded by $\varepsilon_{\mathrm{est}}>0$. For PASS, the in-waveguide channel depends solely on the positions of the pinching antennas and thus can be calculated accurately with negligible error. Channel estimation errors for conventional MIMO and massive MIMO systems can be similarly modeled using (71). The results in Fig. 12 is obtained by optimizing the system with the estimated channel $\hat{\mathbf{h}}_{k}$ and evaluating the performance over the actual channel $\mathbf{h}_{k}$. As expected, increasing channel error bound $\varepsilon_{\mathrm{est}}$ degrades SINR, however, PASS consistently achieves the highest SINR, demonstrating its robustness to channel estimation errors.
@@ V Conclusions
However, several key challenges remain unresolved. For instance, the proposed physics model has been applied only to the downlink scenario, whereas modeling for uplink scenarios may be more complex and requires further investigation. Additionally, this study considered the simplest transmit beamforming architecture, where each waveguide is connected to a single RF chain. Exploring hybrid analog-digital beamforming architectures with pinching beamforming, which support fully connected configurations and a flexible number of RF chains, is another promising direction for future research. Finally, practical indoor or even outdoor deployment is also essential to fully realize the benefits of PASS.